A portfolio is priced to return 10% in the upcoming year. The portfolio is composed of the following three stocks:

Stock

Beta

Share Price

# of shares bought

Alcoa

0.8

$20

100

Budweiser

1.2

$40

100

Cisco

????

$20

100

The current risk-free rate in the economy is 2%, while the market portfolio risk premium is 8%. What must the Beta of Cisco be for the portfolio to be fairly priced?

Question 26 options:

0.6

0.8

1.0

1.4

2.13
Answer #1

Alcoa value = share price × value

= 20 × 100

= 2000

Budweiser value = 40 × 100

= 4000

Cisco = 20 × 100

= 2000

Rate of return = risk free return + beta × ( market risk premium)

Alcoa return = 2% + 0.8 × 8%

= 8.4%

Budweiser return = 2% + 1.2 × 8%

= 11.6%

So, return on alcoa share = Rate of return × Value

= 2000 × 8.4%

= 168

Budweiser return = 4000 × 11.6%

= 464

We have given that total return is 10% and the amount invested is ( 2000 + 4000 + 2000) which is 8000.

So, return on portfolio = 8000 × 10%

= 800

Return on cisco share = 800 - 168 - 464

= 168

So, return in percentage of Cisco share = 168 / 2000

= 8.4%

Beta of cisco share calculation

8.4% = 2% + b ( 8%)

b = 6.4% / 8%

b = 0.80

So, the correct answer is 0.80